OpenStudy (anonymous):

Simplify: (sin Θ − cos Θ)2 − (sin Θ - cos Θ)2

OpenStudy (welshfella):

are you sure you have the question right? as its written the answer is 0

OpenStudy (anonymous):

0

OpenStudy (welshfella):

maybe the sign in one of the brackets is '+' ?

OpenStudy (anonymous):

yeah sorry I just copied and pasted, its supposed to read \[(\sin \theta - \cos \theta)^2-(\sin \theta - \cos \theta)^2\]

OpenStudy (anonymous):

its all subtraction

OpenStudy (welshfella):

well the answer is zero because the two parts are the same

OpenStudy (anonymous):

okay thank you

OpenStudy (welshfella):

yw

OpenStudy (tkhunny):

Generally: Stuff - Stuff = 0

OpenStudy (anonymous):

well that makes sense ^